OpenStudy (anonymous):

Factor x^2 + 2root3x + 3

OpenStudy (anonymous):

\[x^2+2\sqrt{3x}+3?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

It just needs to be factored it comes from getting the square of 9x^2 + 18root3x. I took the nine out and ended up with the equation above.

OpenStudy (anonymous):

\[x^2+2\sqrt{3}\sqrt{x}+3\]

OpenStudy (anonymous):

You can't do much with it.

OpenStudy (anonymous):

That is what I was thinking. It comes from a problem dealing with ellipses

OpenStudy (anonymous):

Is the expression like this: x^2 + 2sqrt(3)x + 3 ? The x is outside or inside the square root?

OpenStudy (anonymous):

The equation is like that. The x is inside the squareroot. Iambatman wrote it out correctly on the first comment.

OpenStudy (anonymous):

did u start with \[9x^{2} + 18\sqrt{3}x + 27 \]

OpenStudy (anonymous):

Then it can't be factored but you can complete its square.

OpenStudy (anonymous):

No. I started out with just the first two terms of the equation you wrote above and had to complete the square for the equation. and that is what I got

OpenStudy (anonymous):

Even completing the square will not work if the x is inside the square root

OpenStudy (anonymous):

Oh? How so?

OpenStudy (anonymous):

it will help if you type your original problem (before doing the completing the square) properly, perhaps draw it or use the formula/equation editor or write it on a piece of paper take a picture of it and attach it here

OpenStudy (anonymous):

Because a complete square is of the form (a + b)^2 = a^2 + 2ab + b^2. That is, you need to have an 'a^2' and a '2ab' term in order to complete the square. Here you have x^2 in the form of a^2 but not 2ab because you have 2sqrt(3x), and it does not have an x without the square root.

OpenStudy (anonymous):

Oh okay. Thank you

OpenStudy (anonymous):

navk i think she meant to write this \[9x ^{2} + 18\sqrt{3} * x\] jess correct me if am wrong

OpenStudy (anonymous):

Yes I did. I just relooked at my problem and the x is not covered by the squareroot sign. Kx2bay you are right

OpenStudy (anonymous):

That greatly simplifies it. Now you can =factor it as follows: x^2 + sqrt(3)x + sqrt(3)x + 3 x(x + sqrt(3)) + sqrt(3)(x + sqrt(2)) (x + sqrt(3))^2

OpenStudy (anonymous):

jess will the answer be this